# Robustness of timelike circular orbit topology against particle spin

### v1

@@ Abstract
Based on a detailed study of the motion of spinning test particles within the Mathisson–Papapetrou–Dixon formalism under the Tulczyjew spin-supplementary condition in static, spherically symmetric spacetimes, we investigate the topological properties of timelike circular orbits (TCOs) for such particles. By constructing an auxiliary potential and an associated vector field on the equatorial plane, we compute the topological winding number $W$ for regions between horizons and outside the outermost horizon in asymptotically flat, anti-de Sitter (AdS), and de Sitter (dS) black hole spacetimes. Our results show that between two neighboring horizons (including the cosmological horizon in the dS case), the topological number is W=1W=-1, indicating the presence of at least one unstable TCO. Outside the outermost horizon, we find $W=0$ for both asymptotically flat and AdS black holes, implying that any TCOs must appear in stable-unstable pairs or be absent. These conclusions are independent of the spin orientation (co-rotating or counter-rotating) of the test particle. The analysis is supported by explicit examples in Schwarzschild, Schwarzschild-AdS, and Schwarzschild-dS spacetimes, confirming the general topological predictions. While the effective potential for spinning particles has been previously studied, the topological approach employed here reveals invariant properties that remain robust even when spin is included, thereby highlighting the fundamental influence of the spacetime structure itself.
@@ I Introduction
Circular orbits in black-hole spacetimes are essential tools for probing strong-field gravity, underpinning key astrophysical phenomena such as accretion disks, black-hole shadows, and gravitational-wave emission from extreme mass-ratio inspirals. Their study provides deep insight into the structure of spacetime and the behavior of matter in extreme environments.
@@
In recent years, topological methods have offered a powerful, model-independent framework for analyzing the existence and stability of circular orbits. These approaches provide a geometric, global perspective on orbital structures that transcends the details of specific metrics or matter content. Cunha, Berti and Herdeiro Cunha:2017qtt pioneered a topological framework to study light rings (LRs) or photon spheres (PSs) of ultra-compact objects, demonstrating that non-degenerate LRs always appear in pairs. They then proceeded to investigate the topological properties of the LRs outside a stationary, axisymmetric black hole Cunha:2020azh , proving that at least one unstable LR exists outside the black hole horizon for each rotation sense. Subsequent work, particularly by Wei et al.  Wei:2020rbh , applied Duan’s topological-current $\phi$-mapping theory Duan:1979ucg ; Duan:1984ws to classify circular orbits in asymptotically flat, AdS and dS black holes, uncovering universal topological invariants associated with photon spheres. Duan’s topological current $\phi$-mapping can also provide new insights into the study of black hole and black hole thermodynamics Wei:2021vdx ; Wei:2022dzw .
@@
Although significant progress has been made in understanding null circular orbits, the case of timelike circular orbits (TCOs) is more intricate, owing to their dependence on orbital parameters such as angular momentum. Earlier studies established that, for neutral particles with fixed angular momentum, TCOs generically arise in stable–unstable pairs Wei:2022mzv ; Ye:2023gmk ; this pairing is topologically enforced and independent of black-hole parameters. Building on this, several further studies have been carried out Shahzad:2024oxb , including the case of charged particles Ye:2024sus ; Song:2025vhw .
In realistic scenarios, however, particles can possess intrinsic spin. The motion of spinning test particles is governed by the Mathisson–Papapetrou–Dixon (MPD) equations Mathisson:1937zz ; Papapetrou:1951pa ; Tulczyjew:1959 ; Taub:1964 ; Madore:1969 ; Dixon:1964 ; Dixon:1965:1 ; Dixon:1970:1 ; Dixon:1970:2 ; Dixon:1973:1 ; Dixon:1974:1 ; Dixon:1979 ; Obukhov:2015eqa ; Semerak:1999qc ; Hackmann:2014tga ; Steinhoff:2009tk , which describe the dynamics of pole–dipole particles in curved spacetime. Spin–curvature coupling induces non-geodesic motion, enriching orbital dynamics and producing spin-dependent shifts in orbital energies and radii. The circular orbits of a spinning test particle in different spacetimes have been studied Jefremov:2015gza ; Toshmatov:2019bda ; Toshmatov:2020wky .
@@
A natural question therefore arises: how does a test particle’s spin affect the topological structure of its circular orbits? Does spin alter the topological invariants that characterise orbital stability? Moreover, are these results universal across different asymptotic behaviours—flat, AdS or dS—and in black hole with multiple horizons? We address these questions by employing a topological approach to study the circular orbits of spinning test particles in static, spherically symmetric spacetimes. Using the Tulczyjew spin-supplementary condition [29], we construct an auxiliary potential and an associated vector field on the equatorial plane. By computing the topological winding number $W$, we classify the existence and stability of TCOs in various regions of black-hole spacetimes.
@@
Our results show that the topological number $W$ is independent of both the magnitude and the orientation of the spin—that is, whether the particle is co-rotating or counter-rotating. Specifically, between two neighbouring horizons we obtain W=1W=-1, indicating the presence of at least one unstable TCO. Outside the outermost horizon of asymptotically flat and AdS spacetimes we find $W=0$, implying that any TCOs must appear in stable–unstable pairs or be absent altogether. These conclusions are corroborated by explicit examples in Schwarzschild, Schwarzschild–AdS and Schwarzschild–dS spacetimes.
This study extends the topological classification of circular orbits to spinning test particles, revealing a robust topological structure that persists even in the presence of spin-induced non-geodesic effects. It underscores the universality of topological methods in gravitational physics and opens new avenues for exploring the dynamics of spinning bodies in more complex spacetimes.
@@
The present paper is organized as follows: In Sec. II, we will give a brief review of the equations of motion of spinning extended test bodies, i.e., MPD equations, and introduce Wei et al.’s topological approach to the study of circular orbits. In Sec. III, we will study the topology of circular orbits for spinning test particles between two neighboring horizons. In Sec. IV, we will study the topology of circular orbits for spinning test particles in asymptotically flat, AdS, and dS black holes. In Sec. V, we summarize and discuss our results. In this work, we use geometrized units, setting $G=c=1$.
@@ II TOPOLOGICAL APPROACH FOR CIRCULAR ORBITS
where $f$ is the function of radial coordinate $r$. The equations of motion for spinning test particles up to the pole-dipole order are given by the MPD equations:
@@
where $D/d\lambda$ is the covariant derivative along the particle’s trajectory with the affine parameter $\lambda$ given by $D/d\lambda\equiv u^{a}\nabla_{a}$ and $R^{a}{}_{bcd}$ is the Riemann tensor. The dynamical 4-momentum and kinematical 4-velocity of the particle are denoted by $p^{a}$ and $u^{a}$, respectively and the anti-symmetric spin tensor is denoted by $S_{ab}$ (with $S_{ab}=-S_{ba}$).
@@
To ensure the mathematical consistency of the derived expressions (e.g., avoiding divergences in $p_{\phi}$, see Eq.(16)) and to maintain the timelike character of the particle’s four-velocity $u^{a}$, see Eq.(8), (thus preserving causality outside the event horizon), we restrict our analysis to cases where the spin magnitude is small enough such that [^4] This condition, which appears in the denominator of key quantities like Eq.(16), helps prevent the spin-curvature coupling from inducing significant deviations that might violate the timelike character of $P^{a}$ and $u^{a}$. While $P^{a}P_{a}$ is timelike by construction from the Tulczyjew condition ($P^{a}P_{a}=-\mathcal{M}^{2}$), the four-velocity $u^{a}u_{a}$ is determined by a more complex relation involving $P^{a}$ and the spin-curvature coupling term. If the spin becomes too large, the denominator $2r-f^{\prime}s^{2}$ can vanish or become negative, leading to divergences or sign changes in momenta components ($P_{t},P_{\phi}$) and potentially resulting in a spacelike or null , implying superluminal motion. This condition is a sufficient (but not always strictly necessary) constraint to avoid these unphysical scenarios and ensure we work within the regime where the pole-dipole approximation and our analysis are valid. While this holds for the spin magnitudes and regions considered, extreme values of $|s|$ (comparable to the black hole mass $M$) or regions very close to the horizon in near-extremal black holes may violate it. Such regimes require a different treatment, which is beyond the scope of this analysis.:
@@
Putting eq.(16) and (17) into eq.(14), we get the result that
@@
in a flat vector space. From Eqs.(21) (24) and (25), we have
@@ III Between two neighboring horizons
We adopt the contour $C$, where $C=\sum_{i}l_{i}$ where $l_{1}\sim l_{4}$$\{r_{\mathrm{out}}=r_{h_{2}},0\leq\theta\leq\pi\}\cup\{\theta=\pi,r_{h_{1}}\leq r<r_{h_{2}}\}\cup\{r_{\mathrm{in}}=r_{h_{1}},0\leq\theta\leq\pi\}\cup\{\theta=0,r_{h_{1}}\leq r<r_{h_{2}}\}$, as illustrated in Fig. 3.
@@
This indicates that for fixed $L$ and $s$, there is always at least one unstable TCO between $r_{h_{1}}$ and $r_{h_{2}}$ Cunha:2020azh ; Wei:2022mzv . This result is independent of the particle’s spin and orientation.
@@ IV.1 Asymptotically flat black holes
$\phi_{l_{1}}^{r}(r\to\infty)\sim\frac{M\mathcal{M}}{r^{2}}$ (42)
@@
This implies that any TCOs with fixed $L$ and $s$ must appear in stable-unstable pairs or be absent. This result is independent of spin.
@@
For a Schwarzschild black hole:
$f(r)=1-\frac{2M}{r}\;.$ (45)
Set $M=1$ and $\mathcal{M}=0.5$. The horizon is located at $r_{h}=2$. The effective potential becomes
@@
    For $L=0.3$ and $1\leq s\leq 1$, the condition $2r-f^{\prime}s^{2}|_{r_{h}}>0$ is satisfied. No TCOs exist (Fig. 6).
    Figure 6: The graph of the potential function (46) for $L=0.3$ and $1\leq s\leq 1$. The red points and the blue boxes represent the maximum values and the minimum values of (46) for a given value of $s$, respectively.
@@
    For $L=1$ and $1\leq s\leq 1$, the condition $2r-f^{\prime}s^{2}|_{r_{h}}>0$ is satisfied. Pairs of stable and unstable TCOs exist (Fig. 7).
    Figure 7: The graph of the potential function (46) for $L=1$ and $1\leq s\leq 1$. The red points and the blue boxes represent the maximum values and the minimum values of (46) for a given value of $s$, respectively.
The numerical evidence from the graphical results for both $L=0.3$ (no TCOs) and $L=1$ (stable-unstable TCO pairs) confirms the topological conclusion $W=0$, which is independent of spin [^5] The angular momentum $L$ must exceed a critical value to provide sufficient centrifugal support for the existence of stable-unstable TCO pairs, otherwise, no TCOs exist..
@@ IV.2 Asymptotically AdS black holes
TCOs must occur in pairs or be absent, independent of spin (as long as $l>|s|$).
@@
Set $M=1$, $\mathcal{M}=0.5$ and $l=10$. The horizon is located at $r_{h}\approx 1.9283$. The behavior of $f(r)$ in Sch-AdS black hole is shown in Fig. 10.
@@
    For $L=0.3$ and $0.25\leq s\leq 0.25$, the conditions $2r-f^{\prime}s^{2}|_{r_{h}}>0$ and $l>|s|$ are satisfied. No TCOs exist (Fig. 11).
    Figure 11: The graph of the potential function (52) for $L=0.3$ and $0.25\leq s\leq 0.25$. The red points and the blue boxes represent the maximum values and the minimum values of (52) for a given value of $s$, respectively.
@@
    For $L=2$ and $1\leq s\leq 1$, the conditions $2r-f^{\prime}s^{2}|_{r_{h}}>0$ and $l>|s|$ are satisfied. Pairs of stable and unstable TCOs exist (Fig. 12).
    Figure 12: The graph of the potential function (52) for $L=2$ and $1\leq s\leq 1$. The red points and the blue boxes represent the maximum values and the minimum values of (52) for a given value of $s$, respectively.
The numerical evidence from the graphical results for both $L=0.3$ (no TCOs) and $L=2$ (stable-unstable TCO pairs) confirms the topological conclusion $W=0$, which is independent of spin.

### v2

# Robust topological invariants of timelike circular orbits for spinning test particles in black hole spacetimes

@@ Abstract
The spin-curvature coupling in the Mathisson–Papapetrou–Dixon (MPD) formalism induces non-geodesic motion, shifting the orbital parameters of spinning test particles in black hole spacetimes. We investigate whether these quantitative shifts alter the qualitative, global structure of the orbit manifold. Using a topological approach, we study timelike circular orbits (TCOs) for spinning particles in static, spherically symmetric spacetimes. By constructing an auxiliary vector field, we compute the topological winding number $W$ in horizon-bounded regions of asymptotically flat, anti–de Sitter (AdS), and de Sitter (dS) backgrounds. We find that $W$ is robust against both the magnitude and direction of the particle’s spin: between two horizons, W=1W=-1, guaranteeing at least one unstable TCO; outside the outermost horizon in asymptotically flat and AdS spacetimes, $W=0$, enforcing that TCOs must appear in stable–unstable pairs or be absent. This spin independence reveals that the fundamental orbital structure is a property of spacetime geometry itself, not of the particle’s spin. We validate this with quantitative examples in Schwarzschild, Schwarzschild-AdS, and Schwarzschild-dS spacetimes, showing explicit spin-induced TCO shifts while confirming the invariant topology. This result provides a topological foundation for interpreting gravitational waveforms from extreme mass-ratio inspirals involving spinning secondaries.
@@ I Introduction
Circular orbits in black-hole spacetimes are essential tools for probing strong-field gravity, underpinning key astrophysical phenomena such as accretion disks, black-hole shadows, and gravitational-wave emission from extreme mass-ratio inspirals (EMRIs). Their study provides deep insight into the structure of spacetime and the behavior of matter in extreme environments.
@@
In recent years, topological methods have offered a powerful, model-independent framework for analyzing the existence and stability of circular orbits. These approaches provide a geometric, global perspective on orbital structures that transcends the details of specific metrics or matter content. Cunha, Berti, and Herdeiro Cunha:2017qtt pioneered a topological framework to study light rings (LRs) or photon spheres (PSs) of ultra-compact objects, demonstrating that non-degenerate LRs always appear in pairs. They then proceeded to investigate the topological properties of the LRs outside a stationary, axisymmetric black hole Cunha:2020azh , proving that at least one unstable LR exists outside the black hole horizon for each rotation sense. Subsequent work, particularly by Wei et al. Wei:2020rbh , applied Duan’s topological-current $\phi$-mapping theory Duan:1979ucg ; Duan:1984ws to classify circular orbits in asymptotically flat, AdS, and dS black holes, uncovering universal topological invariants associated with photon spheres. Duan’s topological current $\phi$-mapping can also provide new insights into the study of black hole thermodynamics Wei:2021vdx ; Wei:2022dzw .
@@
Although significant progress has been made in understanding null circular orbits, the case of timelike circular orbits (TCOs) is more intricate, owing to their dependence on orbital parameters such as angular momentum. Earlier studies established that, for neutral particles with fixed angular momentum, TCOs generically arise in stable–unstable pairs Wei:2022mzv ; Ye:2023gmk ; this pairing is topologically enforced and independent of black-hole parameters. Building on this, several further studies have been carried out Shahzad:2024oxb , including the case of charged particles. It has been shown that the electric charge of the particle indeed alters the topological properties of its circular orbits Ye:2024sus ; Song:2025vhw .
However, this framework is incomplete for describing realistic astrophysical systems. In EMRIs targeted by future gravitational wave detectors like LISA, the secondary object could be a spinning neutron star or black hole. Similarly, accretion disk dynamics and the motion of dark matter candidates with non-zero spin involve particles where spin-curvature coupling becomes significant. The motion of spinning test particles is governed by the Mathisson–Papapetrou–Dixon (MPD) equations Mathisson:1937zz ; Papapetrou:1951pa ; Tulczyjew:1959 ; Taub:1964 ; Madore:1969 ; Dixon:1964 ; Dixon:1965:1 ; Dixon:1970:1 ; Dixon:1970:2 ; Dixon:1973:1 ; Dixon:1974:1 ; Dixon:1979 ; Obukhov:2015eqa ; Semerak:1999qc ; Hackmann:2014tga ; Steinhoff:2009tk , which extend geodesic motion to include dipole interactions with spacetime curvature. This coupling introduces quantitative shifts in orbital energies and radii Jefremov:2015gza ; Toshmatov:2019bda ; Toshmatov:2020wky , but its effect on the global topological structure of the orbit space remains unknown.
This gap in understanding motivates our central research question: does the particle’s spin, like its charge, alter the topological properties of circular orbits? Although the behavior of effective potentials for spinning particles and the existence of circular orbits have been studied since the pioneering work of Tod, de Felice, and Calvani Tod:1976ud , nd spin is known to produce quantitative shifts in orbital parameters Toshmatov:2019bda ; Toshmatov:2020wky , it remains unknown whether it changes the fundamental topological invariants that characterize the global orbital structure. If these invariants are robust against spin, it would imply a deeper level of universality in orbital dynamics, dictated purely by spacetime geometry.
Understanding whether the topological structure of orbits is robust against spin is crucial for gravitational waveform modeling. If the topology depends on spin, then waveform templates must account for spin-induced topological transitions. Conversely, if it is spin-independent, then certain qualitative features of the inspiral are universal, simplifying template construction for detectors like LISA.
@@
We address this gap by employing a topological approach to study the circular orbits of spinning test particles in static, spherically symmetric spacetimes. The central objective of this work is to determine whether the topological winding number $W$, which classifies the existence and stability of TCOs, remains invariant under the introduction of particle spin. Using the Tulczyjew spin-supplementary condition Tulczyjew:1959 , we construct an auxiliary vector field whose zeros correspond to circular orbits. By computing $W$ in various regions of black-hole spacetimes, we demonstrate that the topological structure is indeed independent of both the spin magnitude and orientation. This represents a significant generalization of previous topological results and reveals a new layer of robustness in orbital dynamics.
The topological approach employed here offers distinct advantages over traditional potential analysis, particularly for spinning particles where the effective potential becomes increasingly complex and parameter-dependent. While previous studies have analyzed specific orbital configurations through numerical scanning of potential extrema, our method provides a global, parameter-independent classification. By constructing an auxiliary vector field whose zeros correspond to circular orbits, we extract topological invariants that characterize the existence and stability of orbits without requiring explicit solutions to the equations of motion. This approach reveals universal patterns that persist despite the nonlinearities introduced by spin-curvature coupling.
@@
Our results show that the topological number $W$ is independent of both the magnitude and the orientation of the spin—that is, whether the particle is co-rotating or counter-rotating. Specifically, between two neighbouring horizons we obtain W=1W=-1, indicating the presence of at least one unstable TCO. Outside the outermost horizon of asymptotically flat and AdS spacetimes we find $W=0$, implying that any TCOs must appear in stable–unstable pairs or be absent altogether. These conclusions are corroborated by explicit examples in Schwarzschild, Schwarzschild-AdS, and Schwarzschild-dS spacetimes.
This study extends the topological classification of circular orbits to spinning test particles, revealing a robust topological structure that persists even in the presence of spin-induced non-geodesic effects. The invariance of topological winding numbers under spin perturbations represents a significant finding: although spin quantitatively shifts orbital radii and energies, it does not alter the fundamental topological structure of the orbit space. This robustness has important implications for gravitational waveform modeling, suggesting that certain topological features of inspiral trajectories may be more universal than previously appreciated.
@@
The present paper is organized as follows: In Sec. II, we give a brief review of the equations of motion of spinning extended test bodies, i.e., MPD equations, and introduce Wei et al.’s topological approach to the study of circular orbits. In Sec. III, we study the topology of circular orbits for spinning test particles between two neighboring horizons. In Sec. IV, we study the topology of circular orbits for spinning test particles in asymptotically flat, AdS, and dS black holes. In Sec. V, we summarize and discuss our results. In this work, we use geometrized units, setting $G=c=1$.
@@ II TOPOLOGICAL APPROACH FOR CIRCULAR ORBITS
where $f$ is a function of the radial coordinate $r$. The equations of motion for spinning test particles up to the pole-dipole order are given by the MPD equations:
@@
where $D/d\lambda$ is the covariant derivative along the particle’s trajectory with the affine parameter $\lambda$ given by $D/d\lambda\equiv u^{a}\nabla_{a}$, and $R^{a}_{bcd}$ is the Riemann tensor. The dynamical 4-momentum and kinematical 4-velocity of the particle are denoted by $P^{a}$ and $u^{a}$, respectively, and the anti-symmetric spin tensor is denoted by $S_{ab}$ (with $S_{ab}=-S_{ba}$).
@@
To ensure the mathematical consistency of the derived expressions (e.g., avoiding divergences in $p_{\phi}$, see Eq.(16)) and to maintain the timelike character of the particle’s four-velocity $u^{a}$, see Eq.(8), (thus preserving causality outside the event horizon), we restrict our analysis to cases where the spin magnitude is small enough such that:
@@
This constraint is physically reasonable for test particles with spin magnitudes small compared to the black hole mass, which is the regime where the pole-dipole approximation remains valid. Putting eq.(16) and (17) into eq.(14), we get the result that
@@
in a flat vector space. From Eqs.(21), (24) and (25), we have
@@ III Between two neighboring horizons
We adopt the contour $C=\sum_{i}l_{i}$ where $l_{1}\sim l_{4}$$\{r_{\mathrm{out}}=r_{h_{2}},0\leq\theta\leq\pi\}\cup\{\theta=\pi,r_{h_{1}}\leq r<r_{h_{2}}\}\cup\{r_{\mathrm{in}}=r_{h_{1}},0\leq\theta\leq\pi\}\cup\{\theta=0,r_{h_{1}}\leq r<r_{h_{2}}\}$, as illustrated in Fig. 3.
@@
This indicates that for fixed $L$ and $s$, there is always at least one unstable TCO between $r_{h_{1}}$ and $r_{h_{2}}$ Cunha:2020azh ; Wei:2022mzv . Crucially, this result is independent of the particle’s spin magnitude $s$ and orientation (co- or counter-rotating). The topological guarantee of an unstable TCO between horizons has physical consequences for accretion processes and orbital dynamics in multi-horizon spacetimes. The presence of such orbits may influence the stability of matter configurations and energy extraction mechanisms in these regions. Our analysis extends the topological guarantee of unstable TCOs, previously known for non-spinning particles, to the general case of spinning bodies, revealing the geometric origin of this feature.
@@ IV.1 Asymptotically flat black holes
$\phi_{l_{1}}^{r}(r\to\infty)\sim\frac{M\mathcal{M}}{r^{2}}\;,$ (42)
@@
This implies that, for fixed $L$ and $s$, TCOs must either emerge as a stable–unstable pair or be absent altogether, independent of the particle’s spin.
@@
For Schwarzschild black hole
$f(r)=1-\frac{2M}{r}\;,$ (45)
the effective potential becomes
@@
Set $M=1$, $\mathcal{M}=0.5$ and $1\leq s\leq 1$. The horizon is located at $r_{h}=2$. The angular momentum $L$ must exceed a critical value to provide sufficient centrifugal support for the existence of stable-unstable TCO pairs, otherwise, no TCOs exist. The critical angular momentum $L_{\mathrm{crit}}$ corresponds to the marginally stable circular orbit (MSCO) Wei:2022mzv . In Schwarzschild black hole, The MSCO is the innermost stable circular orbit (ISCO), and from the condition $V^{\prime}(r)=V^{\prime\prime}(r)=0$, one can easily get $L_{\mathrm{ISCO}}\approx 1.732$ for $s=0$.
@@
    The presence of $s$ significantly affects the TCOs, such as ISCO, as illustrated in Fig. 6. For $s>0$ (co-rotating), the ISCO bifurcates into a stable–unstable pair of TCOs, whereas for $s<0$ (counter-rotating), the TCOs vanish entirely [^4] It is important to emphasize that the ISCO itself does not vanish; rather, its position shifts with the spin parameter $s$. Meanwhile, the specific angular momentum originally associated with the ISCO now corresponds either to a stable–unstable TCO pair or to a complete absence of TCOs.. This indicates that the parameter $s$ modifies the position of the ISCO. Moreover, depending on the magnitude of $s$, this significantly influence can extend to different radial widths regions in the vicinity of the ISCO. For $1\leq s\leq 1$, the corresponding range of $L$ is approximately $1.444\lesssim L\lesssim 1.848$.
    Figure 6: The graph of the potential function (46) for $L=1.732$. The red dots and blue squares represent the local maxima and minima of (46) for a given value of $s$, respectively. The solid red line corresponds to the locus where $s=0$, which acts as the boundary separating the region without TCOs from the region where such orbits emerge in stable–unstable pairs.
@@
    For $L\lesssim 1.444$, no TCOs exist, regardless of the value of $s$ (see Fig. 7).
    Figure 7: The graph of the potential function (46) for $L=1$. The red points and the blue boxes represent the maximum values and the minimum values of (46) for a given value of $s$, respectively.
  • (3).
    For $L\gtrsim 1.848$, pairs of stable and unstable TCOs exist, independent of the value of $s$ (see Fig. 8).
    Figure 8: The graph of the potential function (46) for $L=2$. The red points and the blue boxes represent the maximum values and the minimum values of (46) for a given value of $s$, respectively.
The numerical evidence presented above confirms the topological conclusion that the winding number $W=0$, a result that is independent of the spin parameter. Although the presence of $s$ affects the behavior of TCOs, it does not modify their global topological properties.
@@ IV.2 Asymptotically AdS black holes
This implies that, for fixed $L$ and $s$, TCOs must either emerge as a stable–unstable pair or be absent altogether, irrespective of the particle’s spin, provided $l>|s|$.
@@
Set $M=1$, $\mathcal{M}=0.5$, $l=10$ and $1\leq s\leq 1$. The horizon is located at $r_{h}\approx 1.9283$. The behavior of $f(r)$ in Sch-AdS black hole is shown in Fig. 11.
@@
    In Sch-AdS black hole, $L_{\mathrm{ISCO}}\approx 2.526$ for $s=0$. The presence of $s$ significantly affects the TCOs, such as ISCO (see Fig. 12). For $1\leq s<0$ (counter-rotating) and $0.5\lesssim s<1$ (co-rotating), the ISCO bifurcates into a stable–unstable pair of TCOs, whereas for $0<s\lesssim 0.5$ (counter-rotating), the TCOs vanish entirely. This indicates that the parameter $s$ modifies the position of the ISCO. Moreover, depending on the magnitude of $s$, this significantly influence can extend to different radial widths regions in the vicinity of the ISCO. For $1\leq s\leq 1$, the corresponding range of $L$ is approximately $2.11\lesssim L\lesssim 2.54$.
    Figure 12: The graph of the potential function (52) for $L\approx 2.526$. The red points and the blue boxes represent the maximum values and the minimum values of (52) for a given value of $s$, respectively. The red and the blue line represent $s=0$ and $s\approx 0.5$, respectively.
@@
    For $L\lesssim 2.11$, no TCOs exist, independent of the value of $s$ (see Fig. 13).
    Figure 13: The graph of the potential function (52) for $L=1.5$. The red points and the blue boxes represent the maximum values and the minimum values of (52) for a given value of $s$, respectively.
  • (3).
    For $L\gtrsim 3$, pairs of stable and unstable TCOs exist, independent of the value of $s$ (see Fig. 14).
    Figure 14: The graph of the potential function (52) for $L=3$. The red points and the blue boxes represent the maximum values and the minimum values of (52) for a given value of $s$, respectively.
The numerical evidence presented above confirms the topological conclusion that the winding number $W=0$, a result that is independent of the spin parameter. Although the presence of $s$ affects the behavior of TCOs, it does not modify their global topological properties.